# MM-LLMs: Recent Advances in MultiModal Large Language Models

### v4

@@ 7 Conclusion
In this paper, we embark on a comprehensive exploration of the current MM-LLMs landscape, presenting a synthesis from diverse perspectives enriched by our insights. Acknowledging the dynamic nature of this field, it is plausible that certain aspects may have eluded our scrutiny, and recent advances might not be entirely encapsulated. To tackle this inherent challenge, we’ve established a dedicated website for real-time tracking, using crowdsourcing to capture the latest advancements. Our goal is for this platform to evolve into a continuous source of contributions propelling ongoing development in the field. Given the constraints of page limits, we are unable to delve into all technical details and have provided concise overviews of the core contributions of mainstream MM-LLMs. Looking ahead, we commit to vigilant monitoring and continual enhancement of relevant details on our website, incorporating fresh insights as they emerge.
@@ Appendix A Related Surveys
We have established an open-source website for MM-LLMs researchers, supporting crowdsourced updates and aiming to facilitate collaboration in the MM-LLMs field. We anticipate that this survey will illuminate future research in the MM-LLMs domain.
@@ Appendix C Mainstream PEFT Methods
In addition to the aforementioned PEFT methods, there are several others, including P-tuning Liu et al. (2022), P-tuning v2 Liu et al. (2021a), AdaptBias Fu et al. (2022), Compacter Karimi Mahabadi et al. (2021), AdapterFormer Chen et al. (2022a), XTuner Contributors (2023), P-LoRA Dong et al. (2024b), MoLE Chen et al. (2024), and Delta-LoRA Zi et al. (2023).
@@ Appendix E SOTA MM-LLMs
| RTVLM Li et al. (2024a) | RLHF | I+TT | New question-image pairs based on publicly available images or originally diffusion-generated images Gallegos et al. (2023) | Auto.+Manu. | ✗ | –/–/– | – | 5K |
@@
| MMViG Yan et al. (2024) | RLHF | I+TT | MS-COCO | Manu. | ✗ | 16K/–/– | – | 16K |

### v5

@@ 7 Conclusion
Social Impact
MM-LLMs have the potential to impact society. They can enhance accessibility for individuals with disabilities by improved voice recognition and visual aids, fostering equal access to information. In education, MM-LLMs can revolutionize learning with more interactive experiences, catering to diverse learning styles. In media, they can create more engaging content, enriching the consumer experience. However, the widespread adoption of MM-LLMs also poses risks. Privacy concerns arise from the vast training data, raising issues of data security and user privacy. There is also a risk of exacerbating biases in AI algorithms, as biases in training data can lead to biased outputs. Additionally, the automation of tasks traditionally performed by humans could lead to job displacement, necessitating proactive measures to mitigate potential negative impacts on employment. Overall, while MM-LLMs offer promising opportunities, it is essential to address these challenges to ensure their responsible and equitable deployment.
Acknowledgments
We express our gratitude to the anonymous reviewers for their valuable and insightful comments. This work was supported by JSPS KAKENHI Grant Number JP23K28144.
@@ Acknowledgments
In this paper, we embark on a comprehensive exploration of the current MM-LLMs landscape, presenting a synthesis from diverse perspectives enriched by our insights. Acknowledging the dynamic nature of this field, it is plausible that certain aspects may have eluded our scrutiny, and recent advances might not be entirely encapsulated. To tackle this inherent challenge, we’ve established a dedicated website for real-time tracking, using crowdsourcing to capture the latest advancements. Our goal is for this platform to evolve into a continuous source of contributions propelling ongoing development in the field. Given the constraints of page limits, we are unable to delve into all technical details and have provided concise overviews of the core contributions of mainstream MM-LLMs. Looking ahead, we commit to vigilant monitoring and continual enhancement of relevant details on our website (https://mm-llms.github.io), incorporating fresh insights as they emerge.
@@ Appendix A Related Surveys
We have established an open-source website (https://mm-llms.github.io) for MM-LLMs researchers, supporting crowdsourced updates and aiming to facilitate collaboration in the MM-LLMs field. We anticipate that this survey will illuminate future research in the MM-LLMs domain.
@@ Appendix C Mainstream PEFT Methods
In addition to the aforementioned PEFT methods, there are several others, including P-tuning Liu et al. (2022), P-tuning v2 Liu et al. (2021a), AdaptBias Fu et al. (2022), Compacter Karimi Mahabadi et al. (2021), AdapterFormer Chen et al. (2022a), XTuner Contributors (2023), P-LoRA Dong et al. (2024c), MoLE Chen et al. (2024), and Delta-LoRA Zi et al. (2023).
@@ Appendix E SOTA MM-LLMs
| SoM-LLaVA’s IT Yan et al. (2024a) | SFT | I+TT | ShareGPT4V’s IT, LLaVA-1.5’s IT, CogVLM’s IT | Auto.+Manu. | ✔ | –/–/– | 5 | 695K |
@@
| RTVLM Li et al. (2024c) | RLHF | I+TT | New question-image pairs based on publicly available images or originally diffusion-generated images Gallegos et al. (2023) | Auto.+Manu. | ✗ | –/–/– | – | 5K |
@@
| MMViG Yan et al. (2024b) | RLHF | I+TT | MS-COCO | Manu. | ✗ | 16K/–/– | – | 16K |